OpenStudy (anonymous):

hey guys i need help with a derivative problem. Using definition of the derivative get f'(x) by using f(x)= 2x^2-x+5

OpenStudy (anonymous):

Using the definition of a derivative: \[\lim_{h\to0}{\frac{2(x+h)^2-(x+h)+5-(2x^2-x+5)}{h}}\] \[\lim_{h\to0}{\frac{2(x^2+2xh+h^2-(x+h)+5-(2x^2-x+5)}{h}}\] \[\lim_{h\to0}{\frac{4xh+h^2-h}{h}}\] \[\lim_{h\to0}{4x+h-1\] \[4x-1\]

OpenStudy (anonymous):

\[\lim_{h\to0}{4x+h-1}\] Ooops. That second to last line should be the above.

OpenStudy (anonymous):

i dnt know what grade r u in

OpenStudy (anonymous):

I'm not in a grade...